the perimeter of the quadrilateral is 130 inches. Find the lengths of the sides. The given sides are x+2, 2x+1, x-5, x^2-3x.
Answer by windsolace(8)

A quadrilateral has 4 sides as suggested in the name (quad). Add up the 4 sides given in the question.
(x+2) + (2x+1) + (x-5) + (x^2-3x)
It is given in the question that the perimeter is 130 inches. Hence:
(x+2) + (2x+1) + (x-5) + (x^2-3x) = 130
Simplify the equation:
x^2 - x - 2 = 130
x^2 - x - 132 = 0
By factorisation:
(x - 12)( x + 11)= 0
Therefore x = 12 or x = - 11 (reject since x cannot be negative)
Substitute x = 12 into the 4 sides given:
x + 2 = 12 + 2 = 14
2x + 1 = 2(12) + 1 = 25
x - 5 = 12 - 5 = 7
x^2 - 3x = (12)^2 - 3(12) = 108
Hence, the 4 sides of the quadrilateral are 14 inches, 25 inches , 7 inches and 108 inches.